# A Nutrition Multimodal Photoplethysmography Language Model

### v1

@@ Abstract
Hunger and satiety dynamics shape dietary behaviors and metabolic health, yet remain difficult to capture in everyday settings. We present a Nutrition Photoplethysmography Language Model (NPLM), integrating continuous photoplethysmography (PPG) from wearables with meal descriptions. NPLM projects PPG into embeddings interpretable by language models, enabling joint reasoning over physiology and meal context. Trained on 19,340 participants and 1.1 million meal-PPG pairs, the model improved daily caloric intake prediction by 11% over text-only baselines, with accuracy maintained when 80% of meal text was removed. In an independent validation study (n=140) with controlled dining and detailed meal information, the model replicated these findings. These results demonstrate the value of integrating physiological measurements from consumer wearables with meal information for noninvasive dietary monitoring at scale.
@@ 1 Main
Hunger and satiety regulation are central to metabolism, dietary behavior, and health [perry2012appetite, suzuki2012obesity, rakha2022insights]. However, measuring and predicting such responses in free-living conditions remains a major challenge. Dietary assessment methods rely heavily on self-reported intake, which is burdensome and prone to recall bias [subar2015addressing]. Metabolic monitoring using continuous glucose monitoring (CGM) provides valuable physiological insights but remains invasive, cost-prohibitive, and not designed to directly quantify satiety in healthy individuals [mayer1953glucostatic, bornet2007glycaemic, wyatt2021postprandial]. Moreover, the absence of large-scale datasets linking intake to noninvasive physiological responses has hindered personalized nutrition strategies [min2019food, berry2020human]. We introduce Nutrition PPG Language Model (NPLM), a multimodal foundation model that, given photoplethysmography (PPG) waveforms, can infer meal descriptions and predict hunger–satiety shifts.
Foundation models represent a paradigm shift in medical AI, from task-specific to general-purpose architectures [moor2023foundation]. Large Language Models (LLMs) have demonstrated generalization across complex tasks, including nutrition science [singhal2023large, tu2025towards, mcduff2025towards]. Recent work explores fusing physiological signals with LLMs to enhance activity and health inference [cosentino2024towards], or combining multimodal biosignals for personalized insights [merrill2024transforming, khasentino2025personal]. These approaches leverage strengths of both LLMs and physiological signals for contextually relevant predictions.
@@
Wearable sensors, particularly photoplethysmography (PPG), enable passive capture of physiological signals linked to hunger-satiety regulation. PPG measures microvascular blood-volume dynamics. Such hemodynamic changes are known to occur after ingestion [host1996haemodynamic], making nutrition information potentially reflected in PPG signals. Vascular variables including pulse wave velocity and blood pressure have been observed to change at meal-level timescales following meals with varying macronutrient composition, though the relationship to peripheral arterial properties measurable by wrist-worn devices remains an active area of investigation [fryer2021central, lithander2013postprandial]. PPG-derived features encode demographics, cardiovascular state, metabolic health [large-scale-training], and biological clock [miller2025wearable], adding value beyond behavioral signals [erturk2025beyond]. While PPG has been investigated for numerous applications, this work explores nutrition and satiety—a domain where extracting meaningful insights from biosensor data remains challenging in real-world settings with noise and physiological variability.
@@
NPLM aligns two models: a PPG encoder learning health-rich PPG representations [large-scale-training], and a language model learning text representations [radford2019language]. Given a PPG signal, NPLM learns to generate meal descriptions likely consumed near PPG collection time (Fig. 1(a)). We define “near-meal” as PPG segments from -4 to +4 hours relative to meal time. This temporal window accommodates noisy timestamps in free-living realistic (vs controlled) settings (where apps often discretize meal times to fixed slots like ‘12:00 PM’) while maximizing data availability. Within this window, we select the PPG segment closest to the meal timestamp. Most selected segments fall within 90 minutes of meal time, though the broader window ensures sufficient coverage across participants with varying PPG collection patterns. This objective induces PPG representations predictive of meal content, exposing nutritional signals within passively collected PPG waveforms.
@@
We used data from two studies. First, the Apple Heart & Movement Study (AHMS), a real-world observational study where participants provided informed consent to share health information for research use [truslow2024understanding]. We included 19,34019{,}340 subjects with 1,122,8341{,}122{,}834 self-reported meal descriptions. AHMS was used to pretrain NPLM and predict daily caloric intake changes. Second, the Validation Study, an independent controlled study with 140 participants and 720 meals from a dining setting with detailed meal information. [^1] Growing numbers of restaurants write nutrition information into health apps, making this validation study feasible and applicable to real-world settings. Meal descriptions were automatically recorded from ordering apps, and appetite ratings obtained via post-meal surveys, providing satiety labels at controlled times. We used the Validation Study to predict satiety changes after meals. PPG was passively collected throughout both studies. We did not pre-train NPLM on Validation Study data, demonstrating generalization to this cohort. Data filtering details are in Appendix A.1 and cohort statistics in Table 1.
@@
To quantify alignment success, we define alignment win rate as the proportion of meals where PPG-conditional likelihood exceeds text-only likelihood (see Section 2.5). Using near-meal PPG segments within 4 hours of meal times, we find win rate 0.79 (95% CI 0.78–0.80), indicating PPG reliably encodes meal information and validating our alignment procedure.
As negative controls, we permuted PPG embeddings across subjects, and within subjects. Win rate drops below chance across subjects (0.43, 95% CI 0.42–0.44). It decreases within subjects but remains above chance (0.69, 95% CI 0.67–0.71), suggesting subject-specific physiological signals persisting across time (Figure 1(b), left).
@@
We also aligned meal descriptions to PPG increasingly far away from the meal time (Figure 1(b), right). The alignment strength decreases monotonically with temporal distance from meals, confirming that waveforms closest to meal ingestion carry most relevant dietary information (Fig. 1(b)). PPG signals retain predictive capacity even weeks after meals; individuals exhibit substantial meal diversity (average diversity score 0.54), with weak correlation to performance (r=0.09r=-0.09), indicating long-term signals reflect persistent physiological impacts of dietary patterns rather than memorization.
We evaluated how multimodal representations from NPLM enhance satiety prediction. We investigate two downstream tasks:
  1. 1.
    Daily Energy Intake Prediction (on AHMS) – Predicting whether daily caloric intake exceeds personal mean given PPGs and meal descriptions, serving as a proxy for appetite regulation [drapeau2005appetite, rakha2022insights].
  2. 2.
    Postprandial Satiety Prediction (on Validation Study) – Inferring changes in self-reported hunger-satiety levels following meals, given meal descriptions and PPG captured around meals.
@@
In AHMS, NPLM achieved AUC 0.82 (95% CI 0.814–0.826) for distinguishing days with above-mean caloric intake versus below-mean, compared to text-only baseline 0.74 (95% CI 0.732–0.748) (Fig. 2a). In Validation Study, NPLM without retraining attained AUC 0.71 (95% CI 0.70–0.72) for predicting satiety shifts, versus text-only baseline AUC 0.64 (95% CI 0.63–0.65). This represents 11% relative improvement over text-only baselines across cohorts, with multimodal significantly outperforming unimodal baselines (paired $p<0.0001$).
In Validation Study, NPLM attains slightly lower AUC. We attribute this to distributional shift in text modality: meal descriptions in this study are about 49% shorter (less verbose) than in AHMS. Despite coarser annotation, NPLM generalizes effectively, accommodating varying textual detail. Validation Study provides explicit satiety measurements, whereas AHMS uses daily caloric intake as proxy for appetite regulation [drapeau2005appetite], which may also explain performance differences.
To evaluate NPLM’s resilience to shorter, coarser, or incomplete meal descriptions, we perform ablations on AHMS by shortening or removing tokens, then re-evaluating on daily energy intake. We propose two ablations. First, randomly removing 25%, 50%, and 75% of tokens in meal logs. Second, condensing meal descriptions with LLM-generated summaries (see Section 2.5), producing very short descriptions (e.g., “buttermilk pancakes with blueberries” becomes “pancakes”).
Performance results are in Figure 2. Text-only AUC declined with token removal, and PPG-only provided modest baseline. In contrast, multimodal NPLM retained strong performance: with 50% tokens removed, it outperforms full-text-only baseline. With LLM-condensed descriptions (reducing tokens to 78% of original), NPLM achieved AUC 0.75 (95% CI 0.74–0.76), also outperforming full-text-only baseline. PPG signals partly compensate for missing or abbreviated descriptions, highlighting NPLM’s robustness to sparse or noisy text inputs.
@@
To probe biological plausibility, we train an interpretable surrogate model [sigut2023interpretable] that predicts NPLM’s satiety output from nutrient composition when available. This surrogate model is trained using the available nutrient composition and the predictions from NPLM. We leverage nutrient composition data since previous work has shown that nutrient composition is a predictor of satiety [drapeau2005appetite, tucker2016postprandial] and language models have demonstrated the ability to predict nutrient composition from meal descriptions [dhaliwal2025nutribench]. Looking at the regression coefficients, we observe statistically significant coefficients for fat, carbohydrates, fiber and protein (Figure 2d), consistent with prior work [drapeau2005appetite, tucker2016postprandial, simpson2005obesity, raubenheimer2019protein, hill2000dietary, bornet2007glycaemic]. The relationship between macronutrient composition and satiety is complex and mediated by numerous factors beyond simple nutrient content, including meal timing, energy expenditure, sleep quality, and individual metabolic state. The surrogate model coefficients reflect associations learned by NPLM rather than direct causal mechanisms, and should be interpreted as one simplified view of a multi-factorial process. While we do not directly measure all such factors in this study, prior work has shown that PPG-derived features encode information about behavioral patterns (e.g., activity, sleep), metabolic state, and biological aging [erturk2025beyond, large-scale-training, miller2025wearable]. The here observed benefit of incorporating PPG alongside meal text may reflect the model’s implicit capture, via a single and comprehensive signal, of such contextual factors that mediate satiety dynamics, though we of course cannot disentangle their individual contributions in the current analysis.
@@
Despite promising findings, our study has limitations. Both cohorts rely on self-reported meal logs and appetite ratings introducing noise and bias; we expect stronger results with controlled measurement. Cohort demographics are skewed—AHMS participants are U.S. Apple Watch users, Validation Study from controlled dining—limiting socioeconomic and geographic generalizability. AHMS participants logging meals potentially did so for specific health-related goals (e.g., weight management, diabetes management), which may influence dietary patterns and intake-satiety relationships. Validation Study differed demographically from AHMS, including sex distribution. We did not perform demographic matching between cohorts, as Validation Study was designed to test generalization to an independent population rather than provide a demographically matched replication. While this approach allowed us to assess robustness across different demographic distributions, it limits our ability to make strong inferences about subgroup-specific effects. Methodological limitations include: the $\pm 4$ hour temporal window for PPG-to-meal associations is relatively broad and may include confounders such as additional unlogged intake, physical activity, or stress responses; Validation Study assessed satiety exclusively at midday weekday meals, which may not generalize to other meal times given circadian variation in satiety-related hormones; non-caloric intake (water, coffee) not accounted for despite influencing PPG and satiety; behavioral factors like CGM or GLP-1 agonist use not systematically captured. Lack of randomized or longitudinal intervention data precludes causal inference about whether model-driven feedback could improve outcomes. Real-world deployment would require careful attention to data quality, user adherence, and integration with clinical workflows. Despite these limitations, the multimodal model provides strong measurable and actionable information about nutrition effects in real-world settings.
@@
By aligning wrist PPG with meal text at population scale and validating in an independent controlled cohort, NPLM shows vascular pulse signals from wearables encode reproducible signatures of recent intake and satiety dynamics. This work demonstrates nutrition as one research direction where PPG signals provide physiological insights. Robust performance highlights feasibility of noninvasive, scalable monitoring of hunger–satiety dynamics in real-world settings. These findings motivate prospective studies evaluating clinical benefits and influences on metabolic endpoints (weight management, glycemic control, lipid metabolism) and extending predictions to laboratory biomarkers.
@@
Figure 2: NPLM predicts satiety across multiple experimental conditions. (a) Predicting satiated state across cohorts. We report the area under the receiver operating characteristic curve (AUC) for detecting if subjects are above a personalized satiety threshold across both cohorts, comparing multimodal NPLM against text-only baselines. (b) Relationship between performance and text length in AHMS. Removing text lowers average model performance. Text alone outperforms PPG alone, however, PPG with 50% of the tokens outperforms full text. When removing all text (100%), the results illustrate that PPG alone has limited information, though it is powerful when added to (simplified) meal description. (c) Performance with semantic summarizations of food logs. NPLM with summarized meal descriptions outperforms the original full-text baseline. (d) Surrogate nutrient model coefficients.
@@ 3 Data Availability
Data are not publicly available due to limitations in participant’s consent. Any request for data will be evaluated and responded to in a manner consistent with the specific language in the study protocol and informed consent form. Requests for data should be addressed to one of the corresponding authors (K.V.).
@@ Acknowledgments
Authors wish to thank all participants in the Apple Heart & Movement Study, without whom this research would not be possible, as well as partners at the American Heart Association and Brigham and Women’s Hospital. We also want to thank all participants in the Validation Study, as well as the food service provider. Finally, we would like to thank the following individuals for thoughtful discussion and guidance in support of this work: Salar Abbaspourazad, Matt Bianchi, Justin Dobson, Scott Guelich, Calum MacRae, Udhay Nallasamy, Eduardo Martinez Montes, Rajiv Kumar, Cait Dooley, Jen Block, Prajakta Ranade, Lily Peng, Cameron Olsen, and Saurabh Arora.
@@
All authors are employees of Apple, Inc. and own Apple, Inc. stock. The Apple Heart & Movement Study receives funding from Apple, Inc. and the American Heart Association.

### v2

# Wrist Photoplethysmography Predicts Dietary Information

@@ Abstract
Whether wearable photoplethysmography (PPG) contains dietary information remains unknown. We trained a language model on 1.1M meals to predict meal descriptions from PPG, aligning PPG to text. PPG nontrivially predicts meal content; predictability decreases for PPGs farther from meals. This transfers to dietary tasks: PPG increases AUC by 11% for intake and satiety across held-out and independent cohorts, with gains robust to text degradation. Wearable PPG may enable passive dietary monitoring.
@@ 1 Main
Photoplethysmography (PPG) from consumer wearables have been shown to predict cardiovascular, metabolic, and biological clock information [large-scale-training, miller2025wearable, erturk2025beyond]. Whether PPG signal also captures dietary information—i.e., physiological responses to food intake—remains unknown. We test this hypothesis by aligning PPG representations with meal descriptions using a language model, then measuring whether the alignment carries predictive information about meals. We find that wrist PPG reliably encodes meal-related signals: the signal is strongest near meals, decays with time, persists beyond individual traits, and improves prediction of both caloric intake and satiety in an independent cohort.
@@
Hemodynamic changes following ingestion are well-documented [host1996haemodynamic], and vascular properties including pulse wave velocity and blood pressure change in response to meals with varying macronutrient composition, though the relationship to peripheral arterial properties measurable by wrist-worn devices remains an active area of investigation [fryer2021central, lithander2013postprandial]. PPG measures microvascular blood-volume dynamics, and PPG-derived features can predict demographics, cardiovascular state, and metabolic health [large-scale-training], adding value beyond behavioral signals [erturk2025beyond]. However, whether postprandial signals are detectable in wrist-worn PPG, captured opportunistically in real-world conditions with substantial noise and physiological variability, remains unexplored. We hypothesized that if PPG captures postprandial physiology, conditioning on PPG should improve predictions of meal content.
@@
We introduce Nutrition PPG Language Model (NPLM), which aligns two models: a PPG encoder that learns representations predictive of a variety of health conditions [large-scale-training], and a language model that learns rich text representations [radford2019language]. Text serves as a flexible target modality that can represent meal information at varying levels of detail. Given a PPG signal, NPLM learns to generate meal descriptions likely consumed near PPG collection time (Fig. 1(a)). We define “near-meal” as PPG segments from -4 to ++4 hours relative to meal time. This temporal window accommodates noisy timestamps in free-living settings while maximizing data availability. Within this window, we select the PPG segment closest to the meal timestamp. Most selected segments fall within 90 minutes of meal time. This objective induces PPG representations predictive of meal content, exposing nutritional signals within passively collected PPG waveforms.
@@
We used data from two studies. The Apple Heart & Movement Study (AHMS) was a real-world observational study where participants provided informed consent to share health information for research use [truslow2024understanding]. We included 19,34019{,}340 subjects with 1,122,8341{,}122{,}834 self-reported meal descriptions; AHMS was used to train NPLM and evaluate alignment and downstream intake prediction. The Validation Study, an independent controlled study with 140 participants and 720 meals from a dining facility, provided meal descriptions automatically recorded from ordering apps [^1] Participants ordered meals through restaurant ordering applications, which logged meal contents automatically, reducing self-reporting burden compared to AHMS. and post-meal appetite ratings; we used this cohort to test the ability of our PPG-to-text alignment to generalize and predict satiety without retraining. PPG was passively collected in both studies. Data filtering details are in Appendix A.1 and cohort statistics in Table 1.
@@
To quantify how much meal content information a PPG segment contains, we define alignment win rate as the proportion of meals where PPG-conditional likelihood exceeds the regular text-only likelihood given by a capacity-matched text-only baseline (see Section 2.5). Using near-meal PPG segments within 4 hours of meal times, we find win rate 0.79 (95% CI 0.78–0.80), indicating PPG reliably encodes meal-predictive information.
To confirm this signal reflects genuine meal-PPG relationships, we performed two controls (Figure 1(b), left). First, permuting time-matched PPG across subjects reduced win rate to 0.43 (95% CI 0.42–0.44), below chance (50%), confirming the signal is subject-specific. Second, permuting time-matched PPG within subjects yielded 0.69 (95% CI 0.67–0.71), indicating persistent individual traits. The gap between within-subject permutation (0.69) and true pairing (0.79) demonstrates that PPG carries meal-specific information beyond persistent traits—the signal depends on which meal, not just who the person is.
@@
We also tested temporal specificity by aligning meal descriptions to PPG at increasing distances from meals (Figure 1(b), right). Win rate decreased monotonically with time, confirming that waveforms closest to ingestion carry most dietary information and ruling out static time-of-day confounds. Some predictive capacity persists even weeks after meals; given substantial meal diversity across individuals (average diversity score 0.54) with weak correlation to performance (r=0.09r=-0.09), these long-term signals likely reflect persistent physiological impacts of dietary patterns rather than memorization.
Having established that PPG contains meal-predictive information, we next tested whether this signal improves prediction of diet-related outcomes. We evaluated two downstream tasks: (1) Daily Energy Intake Prediction on AHMS, predicting whether daily caloric intake exceeds personal mean as a proxy for appetite regulation [drapeau2005appetite, rakha2022insights]; and (2) Postprandial Satiety Prediction on the Validation Study, inferring changes in self-reported hunger-satiety levels following meals.
@@
In AHMS, NPLM achieved AUC 0.82 (95% CI 0.814–0.826) for distinguishing days with above-mean caloric intake versus below-mean, compared to text-only baseline 0.74 (95% CI 0.732–0.748), an 11% relative improvement (paired $p<0.0001$; Fig. 2a). In the Validation Study, without retraining, NPLM achieved AUC 0.71 (95% CI 0.70–0.72) for predicting satiety shifts versus text-only baseline 0.64 (95% CI 0.63–0.65; paired $p<0.0001$). The consistent improvement across cohorts, outcomes, and text formats suggests the PPG dietary signal generalizes beyond training conditions. Lower absolute performance in the Validation Study likely reflects distributional shift: meal descriptions were 49% shorter than AHMS, and the outcome (satiety) differs from training (intake).
To evaluate NPLM’s resilience to shorter, coarser, or incomplete meal descriptions, we performed ablations on AHMS by randomly removing words or using an LLM to condense descriptions to single keywords (e.g., “buttermilk pancakes with blueberries” becomes “pancakes”; see Section 2.5). Text-only AUC declined with text removal, and PPG-only provided modest baseline (Fig. 2b-c). In contrast, multimodal NPLM retained strong performance: with 50% of text removed, it still outperformed the full-text-only baseline. Crucially, with LLM-condensed descriptions, NPLM achieved AUC 0.75 (95% CI 0.74–0.76), exceeding full-text-only (0.74). This demonstrates that PPG contributes information beyond what text provides: even minimal meal descriptions combined with PPG outperform detailed text alone, suggesting PPG can compensate for incomplete dietary reporting.
@@
To probe biological plausibility, we trained an interpretable surrogate model [sigut2023interpretable] that predicts NPLM’s dietary predictions from nutrient composition when available. Previous work has shown that nutrient composition predicts satiety [drapeau2005appetite, tucker2016postprandial] and language models can estimate nutrient composition from meal descriptions [dhaliwal2025nutribench]; if NPLM captures satiety-relevant signals, its predictions should correlate with nutrients. We observe statistically significant regression coefficients for fat, carbohydrates, fiber and protein (Figure 2d), consistent with prior work [drapeau2005appetite, tucker2016postprandial, simpson2005obesity, raubenheimer2019protein, hill2000dietary, bornet2007glycaemic]. The relationship between macronutrient composition and satiety is complex and mediated by numerous factors beyond simple nutrient content, including meal timing, energy expenditure, sleep quality, and individual metabolic state. The surrogate model coefficients reflect associations learned by NPLM rather than direct causal mechanisms, and should be interpreted as one simplified view of a multi-factorial process. While we do not directly measure all such factors in this study, prior work has shown that PPG-derived features encode information about behavioral patterns (e.g., activity, sleep), metabolic state, and biological clock [erturk2025beyond, large-scale-training, miller2025wearable]. The observed benefit of incorporating PPG alongside meal text may reflect the model’s implicit capture, via a single and comprehensive signal, of such contextual factors that mediate satiety dynamics.
@@
Despite promising findings, our study has limitations. Both cohorts rely on self-reported meal logs and appetite ratings introducing noise and bias [subar2015addressing]; we expect stronger results with controlled measurement. Cohort demographics are skewed: AHMS participants are U.S. Apple Watch users, Validation Study participants from controlled dining, limiting socioeconomic and geographic generalizability. AHMS participants logging meals potentially did so for specific health-related goals (e.g., weight management, diabetes management), which may influence dietary patterns and intake-satiety relationships. Validation Study differed demographically from AHMS, including sex distribution. We did not perform demographic matching between cohorts, as Validation Study was designed to test generalization to an independent population rather than provide a demographically matched replication. While this approach allowed us to assess robustness across different demographic distributions, it limits our ability to make strong inferences about subgroup-specific effects. Methodological limitations include: the $\pm 4$ hour temporal window for PPG-to-meal associations is relatively broad and may include confounders such as additional unlogged intake, physical activity, or stress responses; Validation Study assessed satiety exclusively at midday weekday meals, which may not generalize to other meal times given circadian variation in satiety-related hormones; non-caloric intake (water, coffee) not accounted for despite influencing PPG and satiety; behavioral factors like CGM or GLP-1 agonist use not systematically captured. Lack of randomized or longitudinal intervention data precludes causal inference about whether model-driven feedback could improve outcomes. Real-world deployment would require careful attention to data quality, user adherence, and integration with clinical workflows. Despite these limitations, the model provides measurable and actionable information about dietary responses in real-world settings.
@@
By aligning wrist PPG with meal descriptions at population scale and validating in an independent cohort, we provide evidence that vascular pulse signals from consumer wearables encode reproducible signatures of recent intake and satiety dynamics. The signal is temporally specific (strongest near meals), meal-modulated (beyond persistent physiological traits), robust to text degradation, and generalizes across cohorts and outcomes. These findings establish nutrition as a domain where noninvasive, scalable PPG provides physiological insights, motivating prospective studies on metabolic endpoints (weight management, glycemic control) and extension to laboratory biomarkers.
@@
Figure 2: NPLM predicts dietary outcomes across multiple experimental conditions. (a) Predicting dietary outcomes across cohorts. We report the area under the receiver operating characteristic curve (AUC) for daily caloric intake classification (AHMS) and postprandial satiety prediction (Validation Study), comparing multimodal NPLM against text-only baselines. (b) Relationship between performance and text length in AHMS. Removing text lowers average model performance. Text alone outperforms PPG alone, however, PPG with 50% of the tokens outperforms full text. When removing all text (100%), the results illustrate that PPG alone has limited information, though it is powerful when added to (simplified) meal description. (c) Performance with semantic summarizations of food logs. NPLM with summarized meal descriptions outperforms the original full-text baseline. (d) Surrogate nutrient model coefficients.
@@ 3 Data Availability
Data are not publicly available. Any request for data will be evaluated and responded to in a manner consistent with the specific language in the study protocol and informed consent form. Requests for data should be addressed to one of the corresponding authors (K.V.).
@@ Acknowledgments
Authors wish to thank all participants in the Apple Heart & Movement Study, without whom this research would not be possible, as well as partners at the American Heart Association and Brigham and Women’s Hospital. We also want to thank all participants in the Validation Study, as well as the food service provider. Finally, we would like to thank the following individuals for thoughtful discussion and guidance in support of this work: Salar Abbaspourazad, Saurabh Arora, Matt Bianchi, Jen Block, Justin Dobson, Cait Dooley, Scott Guelich, Rajiv Kumar, Calum MacRae, Eduardo Martinez Montes, Udhay Nallasamy, Cameron Olsen, Lily Peng, and Prajakta Ranade.
Author Contributions
K.V. conceived the research, designed and conducted the Validation Study, developed the model, and conducted analyses. A.N. contributed to model development and analysis. J.F. and A.C.M. provided methodological guidance. G.S. supervised the research. All authors contributed to writing and approved the final manuscript.
@@
The authors declare no competing non-financial interests but the following competing financial interests: All authors are employees of Apple, Inc. and own Apple, Inc. stock. The Apple Heart & Movement Study received funding from Apple, Inc. and the American Heart Association. The Validation Study was also funded by Apple, Inc.